OpenStudy (anonymous):

10-[3+(2x-1)]=3x PLZ show me show how to solve this.

OpenStudy (anonymous):

first we need to open the bracket.. 10- 3- (2x-1) =3 x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u open the second bracket now?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

10- 3- 2x+1 =3 x does that make sense?

OpenStudy (anonymous):

no how did u do that?

OpenStudy (anonymous):

did u understand how i opened the first bracket??

OpenStudy (anonymous):

no it just looks like u rewrote it without them

OpenStudy (anonymous):

10-[3+(2x-1)]=3x 10- 3- (2x-1) =3 x do u notice the sign change coz thr is a negative sign outside the bracket..

OpenStudy (anonymous):

yah i see the sign change. still confused how & y u changed them.

OpenStudy (anonymous):

hmm...can u do this... 5 - ( 3 + 1) = ??

OpenStudy (anonymous):

5-(3+1) 3+1=5 5-(5)=0 is that right?

OpenStudy (anonymous):

lol no...lemme rephrase it. Find the value of 5-(3+1) .

OpenStudy (anonymous):

What?! ok totaly lost now.

OpenStudy (anonymous):

I think Farha is replying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1) you first have to do everything in the bracket 2) then it will become 10-[2+2x]=3x 3) your next step is to get everything out of the bracket. you see their is a subtraction sign in front of the brackets, so consider it to be -1 and distribute it to everything inside the bracket 4) then your answer will be 10-2-2x=3x 5) then combine like terms to get 8-2x3x 6) add2x to both sides to get 8=5x 7) divide 5 from both sides to get x=5/8 hope that helps :D

OpenStudy (anonymous):

THANK YOU!! That's EXACTLY what i needed.